A ring of radius 2cm carries a uniformly distributed positive total charge of 2microcoulombs . calculate the electric field at a point P, lying 3cm from it center along the central axis perpendicular to the plane of the ring.
1
Expert's answer
2021-09-16T10:25:51-0400

Let’s consider the small charge element "dq" on the ring. The total field "dE" at point "P" is the superposition of all fields due to charge elements around the ring. The electric field "dE" has two components: "dE_x=dEcos\theta" along the "x"-axis and "dE_y=dEsin\theta" along the y-axis. During the integration around the ring the non-"x" components will be cancelled by the symmetry. Thus, we get only the "x"-component:


"dE_x=dEcos\theta=\dfrac{kdq}{r^2}cos\theta."

We can find "cos\theta" from the geometry of the task:


"cos\theta=\dfrac{x}{r}."

Therefore, we can write:


"dE_x=\dfrac{kdq}{r^2}\dfrac{x}{r}."

We can find find "r" from the Pythagorean theorem:


"r=\sqrt{x^2+a^2},"

here, "a" is the radius of the ring.

Finally, we get:


"dE_x=\dfrac{kdq}{(x^2+a^2)}\dfrac{x}{\sqrt{x^2+a^2}}=\dfrac{kx}{(x^2+a^2)^{3/2}}dq."

After integrating, we get:


"E_x= \int\dfrac{kx}{(x^2+a^2)^{3/2}}dq,""E_x= \dfrac{kx}{(x^2+a^2)^{3/2}} \int dq=\dfrac{kx}{(x^2+a^2)^{3/2}}Q."

Finally, we can calculate the electric field at a point "P" lying 3 cm from it center along the central axis perpendicular to the plane of the ring:


"E_x(x=0.03\ m)=\dfrac{8.99\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot 0.03\ m}{((0.03\ m)^2+(0.02\ m)^2)^{3/2}}\cdot2\cdot10^{-6}\ C,""E_x(x=0.03\ m)=1.15\cdot10^7\ \dfrac{N}{C}."